# Object Fidelity Diffusion for Remote Sensing Image Generation

### v2

@@ Introduction
LayoutDiff 37.60 0.015 0.447 70.32 7.01 21.73 0.015 0.288 77.56 21.43
GLIGEN 35.06 0.010 0.622 76.41 6.51 39.79 0.026 0.357 76.19 15.58
AeroGen 27.78 0.013 0.563 81.69 55.38 26.65 0.017 0.298 81.91 44.85
$C$-Diff 49.62 0.024 0.685 82.61 42.17 32.40 0.019 0.279 81.63 49.62
Ours 24.92 0.011 0.312 82.55 58.99 20.84 0.014 0.271 83.79 55.68
Table 1: Quantitative comparison with SOTA methods on DIOR and DOTA. We evaluate performance on generation fidelity (FID, KID, CMMD) and layout consistency (CAS, YOLOScore). OF-Diff demonstrates superior overall performance. The best and second best results are highlighted. Lower is better ($\downarrow$), higher is better ($\uparrow$).
@@
7
mAP mAP50 mAP75 mAP mAP50 mAP75
Baseline 30.51 52.84 32.10 38.09 66.31 38.44
LayoutDiff 29.81 52.14 30.36 38.91 66.75 40.37
GLIGEN 28.48 51.27 29.21 38.84 66.10 40.24
AeroGen 31.53 53.37 33.60 38.45 67.09 39.07
$C$-Diff 31.82 53.48 33.97 38.51 66.52 39.02
Ours 32.71 54.44 34.05 40.03 67.89 42.20
Table 2: Trainability ($\uparrow$) comparison on DIOR and DOTA. ‘Baseline’ denotes accuracy with the unaugmented dataset.
DDPO fine-tuning
To enhance the diversity of the distribution of data generated by the fine-tuned model and maintain better consistency with the distribution of real images (Schulman et al. 2015, 2017), denoising diffusion policy optimization (DDPO) (Black et al. 2023) is applied in the post-training of OF-Diff. DDPO regards the denoising process of the diffusion model as a multi-step Markov decision process (MDP). The mapping relationship is defined as follows:
$\pi(a_{t}\mid s_{t})\triangleq p_{\theta}(\mathbf{x}_{t-1}\mid\mathbf{x}_{t},c)$ (8)
$P(s_{t+1}\mid s_{t},a_{t})\triangleq(\delta_{c},\,\delta_{t-1},\,\delta_{\mathbf{x}_{t-1}})$ (9)
$\rho_{0}(s_{0})\triangleq\bigl{(}p(c),\,\delta_{T},\,\mathcal{N}(0,I)\bigr{)}$ (10)
$R(s_{t},a_{t})\triangleq\begin{cases}r(\mathbf{x}_{0},c),&\text{if }t=0,\\ 0,&\text{otherwise}.\end{cases}$ (11)
where $\delta_{z}$ denotes the Dirac delta distribution whose probability density is zero everywhere except at $z$. The symbols $s_{t}$ and $a_{t}$ represent the state and action at time $t$, respectively. Specifically, $s_{t}$ is defined as the tuple composed of the condition $c$, the time step $t$, and the noisy image $\mathbf{x}_{t}$ at that time, whereas $a_{t}$ is defined as the noisy image $\mathbf{x}_{t-1}$ from the preceding time step. The policy is denoted by $\pi(a_{t}\mid s_{t})$, the transition kernel by $P(s_{t+1}\mid s_{t},a_{t})$, the initial state distribution by $\rho_{0}(s_{0})$, and the reward function by $R(s_{t},a_{t})$. To optimize the policy $\pi(a_{t}\mid s_{t})$ so as to maximize the cumulative reward $\mathbb{E}_{\tau\sim p(,\cdot\mid\pi)}!\Bigl{[}\sum_{t=0}^{T}R(\mathbf{s}_{t},\mathbf{a}_{t})\Bigr{]}$, the gradient $\hat{g}$ is computed as follows:
$\hat{g}$= $\mathbb{E}\!\biggl{[}\sum_{t=0}^{T}\frac{p_{\theta}(\mathbf{x}_{t-1}\mid c,\ t,\ \mathbf{x}_{t})}{p_{\theta^{\prime}}(\mathbf{x}_{t-1}\mid c,\ t,\ \mathbf{x}_{t})}$ (12)
$\cdot r(\mathbf{x}_{0},c)\cdot\nabla_{\theta}\log p_{\theta}(\mathbf{x}_{t-1}\mid c,\ t,\ \mathbf{x}_{t})\biggr{]}$
Figure 5: Qualitative results on DIOR and DOTA. OF-Diff is more realistic and fidelity compared to other methods.
The reward functions based on K-Nearest Neighbor (KNN) and KL divergence are introduced, respectively, to optimize the diversity of generated data and the distribution consistency between generated data and real data. $\omega$ is the weight parameter, and $\mathbf{x}_{0}^{\prime}$ is the real image in the dataset.
Experiments
Experimental Settings
Datasets.
  • DIOR-R (Cheng et al. 2022) is a rotated version of the DIOR benchmark (Li et al. 2020), annotated with oriented bounding boxes, 20 categories. We follow the official split, which divides the dataset into training, validation, and test sets with an approximate ratio of 1:1:2. All experiments are trained on the training set and evaluated on the test set.
  • DOTA-v1.0 (Xia et al. 2018) includes 15 common object categories, with 2,806 images and 188,282 instances. This dataset is more challenging due to the presence of numerous dense scenes and small objects. Following MMRotate (Zhou et al. 2022), each image in the training and test sets is cropped into $512\times 512$ tiles, and those without any valid objects are removed. All experiments are conducted on these processed images.
Implementation Details. We train OF-Diff separately on each dataset (DIOR/DOTA), based on the Stable Diffusion 1.5 (Rombach et al. 2022b) pretrained model. Only the ControlNet and shape feature SD decoder are fine-tuned, while all other modules remain frozen. Training is performed using the AdamW optimizer with a learning rate of 1e-5. The global batch size is set to 64, and training runs for 100 epochs.
Benchmark Methods. We compare our method with state-of-the-art L2I generation models for both remote sensing (AeroGen (Tang et al. 2025), CC-Diff (Zhang et al. 2024)), and natural images (LayoutDiffusion (Zheng et al. 2023), GLIGEN (Li et al. 2023)). For a fair comparison, all models are re-trained using our dataset settings, following their official training details respectively.
  • Generation Fidelity. We use FID (Heusel et al. 2017) and KID (Bińkowski et al. 2018) to assess perceptual quality, along with CMMD (Jayasumana et al. 2024), which measures CLIP feature distances between generated and real images to evaluate layout alignment.
  • Layout Consistency. We report CAS (Ravuri and Vinyals 2019) using a pretrained classifier to assess object recognizability, and YOLOScore by applying a pretrained Oriented R-CNN (Xie et al. 2021) (w/. Swin Transformer backbone (Liu et al. 2021), MMRotate) to generated images for instance-level consistency.
  • Shape Fidelity. To assess the geometric quality of generated instances, we perform pairwise comparisons with ground-truth shapes. Each instance pair is cropped, resized to $\times64$, and converted to edge maps. We compute five metrics: IoU, Dice, Chamfer Distance (CD), Hausdorff Distance (HD), and SSIM (Wang et al. 2004).
@@
Qualitative Results
Diversity Results. Figure 6 shows that the images generated by OF-Diff consistently present plausible textures and realistic object shapes. For instance, airplanes rendered at different orientations maintain coherent semantic relationships with their surrounding environments. Even in small-object scenes (some of which are grayscale remote-sensing images from the DOTA dataset), OF-Diff can still generate visually faithful and geometrically accurate results.
  Method Dataset IoU $\uparrow$ Dice $\uparrow$ CD $\downarrow$ HD $\downarrow$ SSIM $\uparrow$
LayoutDiff 0.0497 0.0908 12.037 25.962 0.1667
GLIGEN 0.055 0.1002 12.257 25.850 0.1652
AeroGen DIOR 0.0855 0.153 8.209 20.314 0.2142
$C$-Diff 0.0891 0.1582 8.0909 20.066 0.1963
Ours 0.1009 0.1763 7.6579 19.459 0.2691
LayoutDiff 0.0402 0.0748 15.229 30.202 0.2194
GLIGEN 0.0645 0.1182 10.432 23.196 0.1967
AeroGen DOTA 0.0863 0.1536 8.1386 20.687 0.2261
$C$-Diff 0.0692 0.1255 9.6226 21.247 0.2171
Ours 0.1205 0.2045 6.6317 17.311 0.2938
@@
YOLO
[…]
LayoutDiff 44.58 0.018 0.539 29.34 10.37 53.07
GLIGEN 39.56 0.013 0.444 66.36 2.13 52.68
AeroGen 28.62 0.013 0.276 80.78 46.36 55.11
$C$-Diff 49.92 0.024 0.513 78.01 51.74 53.72
Ours 24.18 0.012 0.271 83.34 49.59 56.65
Figure 6: The diversity of different results from the same OF-Diff model.
Figure 7: AP50 on DIOR and DOTA.
Quantitative Results
Generation Fidelity and Consistency. We compared OF-Diff with state-of-the-art generation methods in remote sensing, including layoutDiffusion (Zheng et al. 2023), GLIGEN (Li et al. 2023), AeroGen (Tang et al. 2025), and CC-Diff (Zhang et al. 2024). The performance of these methods is reported in Table 1. OF-Diff achieved nearly the best performance in both generation fidelity metrics (FID, KID, CMMD) and layout consistency metrics, especially on the DOTA dataset.
Trainability of Object Detection. Following the data enhancement protocol in (Chen et al. 2023), we double the training samples using OF-Diff and assess detection results with the expanded dataset. As shown in Table 2, OF-Diff performs the best on both DIOR and DOTA. Specifically, OF-Diff improves mAP by 2.2% on DIOR and 1.94% on DOTA compared to baseline. Notably, the performance of several polymorphic and small object classes shows significant improvement. According to Figure 7, the AP50 increases by 8.3%, 7.7%, and 4.0% for airplane, ship, and vehicle on DIOR, and 7.1%, 5.9% and 4.4% for swimming pool, small vehicle, and large vehicle on DOTA.
Object-Shape Fidelity. We measure the morphological similarity between the generated instances and ground truth by calculating the Intersection over Union (IoU), DICE coefficient, Chamfer distance (CD), Hausdorff distance (HD), and Structural Similarity Index (SSIM), based on the Canny Edge Map. As shown in Table 3, the results demonstrate that OF-Diff attains state-of-the-art performance in all evaluation metrics for object-shape fidelity. Specifically, we first convert the rotated bounding box (R-Box) to a horizontal bounding box (H-Box) and crop the instance with a 20% padding to ensure the full object is captured. The cropped patches are then resized to $64\times 64$ pixels, and their shapes are extracted using cv2.Canny. For a detailed qualitative comparison, Figure 9 visualizes the instance patches and their corresponding edge maps from different methods. Each image set is ordered as follows: Ground Truth, OF-Diff, AeroGen, CC-Diff, GLIGEN, and LayoutDiff, demonstrating our method’s superior ability to adhere to object shapes.
Adaptability of Unknown Layout. To better demonstrate the scalability of these methods, we also generate images based on the unknown layouts during the training phase. According to Table 4, for unknown layout, OF-Diff performs well in terms of generation fidelity, layout consistency, and trainability. In downstream tasks, OF-Diff still delivers a 1.54% mAP gain over the second-best method.
The Detailed Results of Downstream. Table 6 and  7 report the average precision (AP) obtained by the competing generative methods over multiple categories in the downstream tasks. From Tables 6, it can be observed that OF-Diff (ours) achieves a clear advantage in several categories. For instance, OF-Diff achieves superior performance on Airplane (71.3%), Golf Field (75.4%) and Ship (70.5%), with improvements of approximately 5% to 10% over the second-best method. For a few other categories, OF-Diff does not deliver the top AP, yet the gap to the best result remains marginal. Table 7 shows that, on the DOTA dataset, OF-Diff obtains the highest AP in roughly half of all categories and still delivers notable gains in categories such as Small-vehicle (68.3%), Ship (84.4%) and Swimming-pool (67.9%).
ESGM DCLoss DDPO CAS $\uparrow$
YOLOScore $\uparrow$
mAP50 $\uparrow$
80.27 41.20 52.13
82.16 55.08 52.76
81.57 46.27 53.14
81.91 47.74 53.21
82.30 57.83 54.31
82.55 58.99 54.44
@@
Table 5: Ablation study: impact of ESGM, DCLoss, and DDPO on semantic consistency (CAS) and downstream trainability (YOLOScore and mAP50).
Figure 8: The influence of caption on the generation of images in terms of being more realistic and more aesthetically pleasing.
Ablation Study
@@
We assessed the impact of different modules on image generation semantic consistency and downstream trainability by OF-Diff in Table 5. We found that the input of captions would affect the fidelity of image generation. Images generated with captions are more in line with semantic consistency and human aesthetics, but the fidelity of these images decreases. This is equivalent to the data distribution deviating from the real dataset and being more inclined towards the data distribution during pre-training. Therefore, the ablation experiments for each module were conducted based on the absence of caption input. The contribution of each module to the enhancement of image generation fidelity is evaluated by incorporating additional components into the diffusion model with dual-branch. DDPO indicates whether to fine-tune the trained diffusion model through reinforcement learning. Results show that Enhanced Shape Generation Module (ESGM), Diffusion Consistency Loss (DCLoss) and the DDPO based on KNN and KL Divergence effectively improve the performance metrics. Notably, ESGM can substantially improve the YOLOScore by over 10%.
Discussion
@@
As shown in Figure 8, the inclusion of additional captions as input has a significant impact on the outcomes of image generation. Specifically, incorporating captions enhances the aesthetic appeal of the generated images, resulting in richer and more visually pleasing color compositions. However, this improvement comes at a cost: similar to CC-Diff, it leads to a deviation of the generated data distribution from that of the original real data. In contrast, when no additional captions are provided as input, although the generated images may appear less aesthetically refined, their data distribution remains closer to that of real images. Furthermore, experimental results demonstrate that this approach yields superior performance in terms of image quality metrics and downstream task effectiveness.
According to the current experimental results, adding the DDPO strategy does not simultaneously outperform previous results on all metrics. Using reinforcement learning strategies can indeed improve the performance of downstream tasks, but it does not necessarily improve the quality of image generation simultaneously. In other words, reinforcement learning strategies can also purposefully improve the quality of image generation, but this may come at the cost of not improving the performance of downstream tasks.
Limitations
@@
The proposed OF-Diff injects object shape masks extracted from the image layout as controllable conditions into the diffusion model, which effectively enhances object fidelity and improves the generation of small objects. However, this also makes the model heavily dependent on the quality of the extracted shape masks. If the obtained masks are already distorted, the resulting generated images will likewise fail to achieve satisfactory quality.
Conclusion
@@
Existing image generation methods struggle to precisely generate dense small objects and those with complex shapes, such as numerous small vehicles and airplanes in remote sensing images. To address this, we introduce OF-Diff, a dual-branch controllable diffusion model with prior shapes extraction and DDPO. During the training phase, we extract the prior shape of objects to enhance controllability and use a dual-branch diffusion with parameter sharing to improve the model’s learning ability for real images. Therefore, in the sampling phase, OF-Diff can generate images with high fidelity without real images as references. Finally, we fine-tune the diffusion by DDPO that combines KNN and KL divergence to make the synthesized images more realistic and consistent. Extensive experiments demonstrate the effectiveness and superiority of OF-Diff in generating small and difficult objects with complex structures and dense scenes in remote sensing.

### v3

@@ 4.2 Qualitative Results
LayoutDiff 37.60 0.015 0.447 70.32 7.01 52.14 21.73 0.015 0.288 77.56 21.43 66.75
GLIGEN 35.06 0.010 0.622 76.41 6.51 51.27 39.79 0.026 0.357 76.19 15.58 66.10
AeroGen 27.78 0.013 0.563 81.69 55.38 53.37 26.65 0.017 0.298 81.91 44.85 67.09
$C$-Diff 49.62 0.024 0.685 82.61 42.17 53.48 32.40 0.019 0.279 81.63 49.62 66.52
Ours 24.92 0.011 0.312 82.55 58.99 54.44 20.84 0.014 0.271 83.79 55.68 67.89
Figure 5: AP50 on DIOR and DOTA.
@@
LayoutDiff 0.0497 0.0908 12.037 25.962 0.1667 0.0402 0.0748 15.229 30.202 0.2194
GLIGEN 0.055 0.1002 12.257 25.850 0.1652 0.0645 0.1182 10.432 23.196 0.1967
AeroGen 0.0855 0.153 8.209 20.314 0.2142 0.0863 0.1536 8.1386 20.687 0.2261
$C$-Diff 0.0891 0.1582 8.0909 20.066 0.1963 0.0692 0.1255 9.6226 21.247 0.2171
Ours 0.1009 0.1763 7.6579 19.459 0.2691 0.1205 0.2045 6.6317 17.311 0.2938
@@
LayoutDiff 44.58 0.018 0.539 29.34 10.37 30.41 53.07 32.07
GLIGEN 39.56 0.013 0.444 66.36 2.13 30.06 52.68 31.29
AeroGen 28.62 0.013 0.276 80.78 46.36 32.98 55.11 34.26
$C$-Diff 49.92 0.024 0.513 78.01 51.74 32.49 53.72 35.39
Ours 24.18 0.012 0.271 83.34 49.59 33.02 56.65 36.17
@@
Table 4: Ablation study: impact of ESGM, Online-distillation $L_{c}$, and DDPO on semantic consistency (CAS) and downstream trainability (YOLOScore and mAP50).
  ESGM $L_{c}$ DDPO FID $\downarrow$ KID $\downarrow$ CMMD $\downarrow$ CAS $\uparrow$ YOLOScore $\uparrow$ mAP50 $\uparrow$
42.59 0.029 0.965 80.27 41.20 52.13
24.87 0.012 0.428 82.16 55.08 52.76
36.25 0.021 0.596 81.57 46.27 53.14
41.26 0.027 0.815 81.06 42.53 53.41
24.98 0.010 0.313 82.30 57.83 54.31
25.78 0.013 0.368 82.37 58.26 54.17
37.98 0.025 0.613 81.91 47.74 53.21
24.92 0.011 0.312 82.55 58.99 54.44
4.3 Quantitative Results
Generation Fidelity and Consistency. We compared OF-Diff with state-of-the-art generation methods in remote sensing, including layoutDiffusion [zheng2023layoutdiffusion], GLIGEN [li2023gligen], AeroGen [tang2025aerogen], and CC-Diff [zhang2024cc]. The performance of these methods is reported in Table 1. OF-Diff achieved nearly the best performance in both generation fidelity metrics (FID, KID, CMMD) and layout consistency metrics, especially on the DOTA dataset. Additional results are available in the appendix 6.3 for the HRSC2016 dataset.
Trainability of Object Detection. Following the data enhancement protocol in [Chen_Xie_Chen_Wang_Hong_Li_Yeung_2023], we double the training samples using OF-Diff and assess detection results with the expanded dataset. As shown in Table 9 in Appendix 6.6, OF-Diff performs the best on both DIOR and DOTA with mAP improved by 2.2% and 1.94% compared to baseline,respectively. Notably, the performance of several polymorphic and small object classes shows significant improvement. According to Figure 5 (a) and (b), the AP50 increases by 8.3%, 7.7%, and 4.0% for airplane, ship, and vehicle on DIOR, and 7.1%, 5.9% and 4.4% for swimming pool, small vehicle, and large vehicle on DOTA.
Object-Shape Fidelity. We measure the morphological similarity between the generated instances and ground truth by calculating the Intersection over Union (IoU), DICE coefficient, Chamfer distance (CD), Hausdorff distance (HD), and Structural Similarity Index (SSIM), based on the Canny Edge Map. As shown in Table 2, the results demonstrate that OF-Diff attains state-of-the-art performance in all evaluation metrics for object-shape fidelity. Specifically, we first convert the rotated bounding box (R-Box) to a horizontal bounding box (H-Box) and crop the instance with a 20% padding to ensure the full object is captured. The cropped patches are then resized to $64\times 64$ pixels, and their shapes are extracted using cv2.Canny. For a detailed qualitative comparison, Figure 11 in Appendix 6.7 visualizes the instance patches and their corresponding edge maps from different methods. Each image set is ordered as follows: Ground Truth, OF-Diff, AeroGen, CC-Diff, GLIGEN, and LayoutDiff, demonstrating our method’s superior ability to adhere to object shapes.
Adaptability of Unknown Layout. To evaluate robustness of these methods, we also generate images based on the unknown layouts during the training phase. According to Table 3, for unknown layout, OF-Diff performs well in terms of generation fidelity, layout consistency, and trainability. In downstream tasks, OF-Diff still delivers a 1.54% mAP gain over the second-best method.
The Detailed Results of Downstream. Table 10 and 11 in Appendix 6.7 report the average precision (AP) obtained by the competing generative methods over multiple categories in the downstream tasks. From Tables 10, it can be observed that OF-Diff (ours) achieves a clear advantage in several categories. For instance, OF-Diff achieves superior performance on Airplane (71.3%), Golf Field (75.4%) and Ship (70.5%), with improvements of approximately 5% to 10% over the second-best method. For a few other categories, OF-Diff does not deliver the top AP, yet the gap to the best result remains marginal. Table 11 shows that, on the DOTA dataset, OF-Diff obtains the highest AP in roughly half of all categories and still delivers notable gains in categories such as Small-vehicle (68.3%), Ship (84.4%) and Swimming-pool (67.9%).
4.4 Ablation Study
@@
We assessed the impact of different modules on image generation semantic consistency and downstream trainability by OF-Diff in Table 4. We found that the images generated with captions are more in line with semantic consistency and human aesthetics, but the fidelity of these images decreases. This is equivalent to the data distribution deviating from the real dataset and being more inclined towards the data distribution during pre-training. We conduct human/GPT assessments and a fine-grained feature analysis in Appendix 6.6, which collectively reveal the nature of this trade-off. Therefore, the ablation experiments for each module were conducted based on the absence of caption input. The contribution of each module to the enhancement of image generation fidelity is evaluated by incorporating additional components into the diffusion model with online-distillation. DDPO indicates whether to fine-tune the trained diffusion model through reinforcement learning. Results show that Enhanced Shape Generation Module (ESGM), Online-Distillation ($L_{c}$) and the DDPO based on KNN and KL Divergence effectively improve the performance metrics. Notably, ESGM can substantially improve the YOLOScore by over 10%. In addition, we vary the weighting coefficient $\lambda$ in the consistency loss (Eq. 7) to assess its impact on mAP and FID. As shown in Figure 5 (c) and (d), both metrics are optimal at $\lambda=1$.
4.5 Discussion
@@
As shown in Figure 10 in Appendix 6.2, the inclusion of additional captions as input has a significant impact on the outcomes of image generation. Specifically, incorporating captions enhances the aesthetic appeal of the generated images, resulting in richer and more visually pleasing color compositions. However, this improvement comes at a cost: similar to CC-Diff, it leads to a deviation of the generated data distribution from that of the original real data. In contrast, when no additional captions are provided as input, although the generated images may appear less aesthetically refined, their data distribution remains closer to that of real images. A user study from both human and GPT-5 in Table 8 have confirmed this. Further analysis of the generated-image distribution and the impact of aesthetics on performance is provided in the Appendix 6.6.
4.6 Limitations
@@
The proposed OF-Diff injects object shape masks extracted from the image layout as controllable conditions into the diffusion model, which effectively enhances object fidelity and improves the generation of small objects. However, this also makes the model heavily dependent on the quality of the extracted shape masks. If the obtained masks are already distorted, the resulting generated images will likewise fail to achieve satisfactory quality. Therefore, in some scenarios, it is necessary to adopt more suitable mask extraction models (e.g., RemoteSAM, SAM2, SAM3) to obtain higher-quality masks. In future work, we will explore stronger object mask extraction and generation strategies to better ensure object-level generation quality while improving the texture details of the synthesized images. We will also investigate more advanced models and training paradigms to further promote the development of remote-sensing image generation.
5 Conclusion
@@
Existing image generation methods struggle to precisely generate dense small objects and those with complex shapes, such as numerous small vehicles and airplanes in remote sensing images. To address this, we introduce OF-Diff, an online-distillation controllable diffusion model with prior shapes extraction. During the training phase, we extract the prior shape of objects to enhance controllability and use a online-distillation diffusion with parameter sharing to improve the model’s learning ability for real images. Therefore, in the sampling phase, OF-Diff can generate images with high fidelity without real images as references. Finally, we fine-tune the diffusion by DDPO that combines KNN and KL divergence to make the synthesized images more realistic and consistent. Extensive experiments demonstrate the effectiveness and superiority of OF-Diff in generating small and difficult objects with complex structures and dense scenes in remote sensing.